OpenStudy (njabulo):

please need help on ......if 5x/x-2 = a/x^2-5x+6 then find a

OpenStudy (anonymous):

5x / x-2 = a / (x-2)(x-3) multiply thru by x-2: 5x = a / (x-3) a =5x(x-3)

OpenStudy (njabulo):

thank you Guru